Circular No. 6229

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1995 S1 The following additional approximate positions have been reported:

1995 UT  α_{2000}  δ_{2000}  m_{1} Observer
Sept.18.50  9^{h}07^{m}  − 1^{o}57′ 6.5 Morris
18.517 9 06.1 − 1 46  6  Machholz

C. S. Morris (Whitaker Peak, CA). 20×80 binoculars. Very approximate position. Comet strongly condensed, coma diameter ∼ 4′; extremely faint tails 70′ and 45′ long in p.a. ∼ 295^{o} and 5^{o}, respectively.

D. E. Machholz (Colfax, CA). Independent discovery with 27×120 binoculars; coma diameter ∼ 10′.

COMET 6P/D’ARREST M. Fomenkova, University of California, San Diego; and M. J. Mumma,

Goddard Space Flight Center, NASA, write: “We obtained infrared images and spectra with the NASA Infrared Telescope Facility (+ NSFCAM) on Aug. 29–Sept. 1 UT. On Sept. 1.5, the fluxes (measured in W m^{−2} μm^{−1}) through J, K, and L filters in a box of size 1′′.2×1′′.2 centered on the nucleus were: J, 6.8 × 10^{−15}; K, 1.3 × 10^{−15}; L, 3.6 × 10^{−15}. On Aug. 31, lowresolution spectra (3.2–3.8 μm) were acquired using the grism and the L filter. The cometary spectrum is significantly redder than an F5 comparison star. The comet shows excess emission above the thermal continuum at 3.4 μm and possibly also at 3.52 μm.”

COMET C/1995 O1 (HALE-BOPP) Fomenkova and Mumma also communicate: “We obtained infrared im-

ages of this comet (details as above) on Sept. 1.3 UT, yielding the following fluxes (in W m^{−2} μm^{−1}) in a box of size 1′′.2 × 1′′.2 centered on the nucleus: J, 3.5 × 10^{−15}; K, 5.6 × 10^{−15}.”

COMET 58P/JACKSON-NEUJMIN Total visual magnitude estimates: Sept. 4.39 UT, 13.5 (C. S. Morris,

Pine Mountain Club, CA, 0.26-m reflector); 16.23, 13.4 (Morris); 17.24, 12.9 (A. Hale, Cloudcroft, NM, 0.41-m reflector).

1995 September 18 Daniel W. E. Green